OpenStudy (anonymous):

Help! How do I do this problem? The cost, C, in millions of dollars of producing T million tons of steel is modeled by the equation C = 220T + 1890. Solve the equation for T. The steel produced is expected to be sold at a price of $310 per ton. Write and solve an equation to find the amount of steel produced if the selling price is equal to the cost of production.

OpenStudy (aum):

C = 220T + 1890 To solve for T, isolate T. Subtract 1890: C - 1890 = 220T divide by 220 (C - 1890) / 220 = T or T = (C - 1890) / 220

OpenStudy (anonymous):

Makes sense! Thank you so much aum!

OpenStudy (aum):

For the second part, C = 220T + 1890 cost of producing one ton is C/T. C/T = 220 + 1890/T = selling price = $310 220 + 1890/T = 310 subtract 220: 1890/T = 310 - 220 = 90 multiply by T: 1890 = 90T divide by 90: 1890/90 = T T = 21 21 million tons of steel.